What is the length, in feet, of the hypotenuse of a right triangle with legs that are 6 feet long and 7 feet long, respectively?: What is the length, in feet, of the hypotenuse of a right triangle with legs that are 6 feet long and 7 feet long, respectively?
Answer by checkley77(3848)

6^2+7^2=x^2
36+49=x^2
85=x^2
x=sqrt85
x=9.22 Answer.